# Almost equivalences between Tamarkin category and Novikov sheaves

### v2

@@ 1 Introduction
In symplectic geometry, there exists a series of extra variables, which are (expected to be) just various aspects of one variable.
@@
    In microlocal sheaf theory, it is called $t$ and introduced by Tamarkin [Tam18], which was also envsioned by Sato in relation with WKB analysis [SK14].
@@
For example, in [Kuwal], we discussed the relationship between (1) and (2). Also, the relationship between (1) and (3) can be seen from the work of Tamarkin, and later clarified by [AI, IK].
@@
In [Kuwal], we found the Novikov ring a posteriori after the construction of the equivariant version of Tamarkin’s category. In this paper, we try to build the Novikov ring a priori into sheaves. Hence, it gives an alternative model of the category of sheaf quantizations.
We would like to formulate our main theorem now. Let $\mathbb{K}$ be a field. Let $M$ be a manifold and $\mathbb{R}_{t}$ be the real line. We denote the discrete additive group of the real numbers by $\mathbb{R}_{d}$. We then have the equivariant derived category $\mathrm{Sh}^{\mathbb{R}_{d}}(M\times\mathbb{R}_{t},\mathbb{K})$ of $\mathbb{K}$-modules sheaves. We quotient this category by the sheaves with non-positive microsupport and denote it by $\mu(T^{*}M,\mathbb{K})$. On the other hand, we have the universal Novikov ring $\Lambda_{0}$ over $\mathbb{K}$. We denote the derived category of $\Lambda_{0}$-module sheaves by $\mathrm{Sh}(M,\Lambda_{0})$.
@@
$\mathfrak{A}\colon\mu(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,\Lambda_{0}).$ (1.1)
The almost image satisfies the derived completeness.
Here the term “almost” comes from the almost mathematics [GR03]. Roughly speaking, the above theorem without “almost” holds after negating almost zero modules. The precise meaning will be explained in the body of the paper. The use of the almost mathematics in symplectic geometry also happened in [Fuk21].
@@
A variant of the main theorem was also described in [Kuw21] where we used the finite Novikov ring $\mathbb{K}[\mathbb{R}_{\geq 0}]$, and compared it with the enhanced ind-sheaves, to reformulate the holonomic Riemann–Hilbert correspondence [DK16].
When we try to relate (1) and (3) categorically, it is easier to construct a functor valued in the sheaves of modules over the Novikov ring. For example, our construction is crucial when we would like to relate sheaf theory and Fukaya category. In the subsequent publication, we plan to combine this result with a generalization of Viterbo’s construction [Vit] to upgrade the Fukaya–sheaf correspondence [GPS20] to an equivalence over the Novikov ring. An integral version of the statement is carried out in a joint work with Petr and Shende [KPS] in a different way, but we use some computations from this article. Also, we plan to rewrite the preprint [Kuw22] using the present formalism.
Aside from the above applications, in this paper,we will explain two applications of (the philosophy) of the theorem.
The first one is a first step toward nonconic microlocal sheaf theory of sheaves valued in modules over real valuation rings. Namely, for a manifold $M$, a real valuation ring $R$, and a sheaf $\mathcal{E}$ of modules over $R$, we can define a subset $\mu\mathrm{supp}(\mathcal{E})\subset T^{*}M$ by interpreting Tamarkin’s non-conic microsupport. This refines Kashiwara–Schapira’s microsupport [KS94].
@@
Let $\mathbb{K}$ be a field.
@@ Acknowledgment
This work is supported by JSPS KAKENHI Grant Numbers 22K13912, 23H01068, and 20H01794. I’d like to thank Yuichi Ike for related discussions. I also thank Tomohiro Asano for pointing out a mistake.
@@ 2.1 Definition
Let $\mathbb{R}$ be the 1-dimensional Euclidean vector space. Let $\mathbb{G}$ be a subgroup of $\mathbb{R}$. Then $\mathbb{R}_{\geq 0}\cap\mathbb{G}$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$. We denote the indeterminate corresponding to $a\in\mathbb{R}_{\geq 0}\cap\mathbb{G}$ by $T^{a}$. Let ||$\cdot$|| ⋅ | be the Euclidean norm of $\mathbb{R}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$ generated by $T^{a}$’s with $a>r$ by $\mathfrak{m}(r)$. Obviously, $\mathfrak{m}(r^{\prime})\supset\mathfrak{m}(r)$ if $r>r^{\prime}$. Hence $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]/\mathfrak{m}(r)$ forms a projective system.
@@
$\Lambda_{0}:=\Lambda_{0}^{\mathbb{R}}:=\lim_{\begin{subarray}{c}\longleftarrow\\ a\rightarrow+\infty\end{subarray}}\mathbb{K}[\mathbb{R}_{\geq 0}]/T^{a}\mathbb{K}[\mathbb{R}_{\geq 0}].$ (2.2)
@@ 2.2 Modules over the Novikov ring
Let $M$ be a topological space. Let $R_{M}$ be the constant sheaf valued in $R$. We denote the abelian category $R_{M}$-modules by $\mathrm{Mod}^{\heartsuit}(R_{M})$. We similarly consider the derived category of $\mathrm{Mod}^{\heartsuit}(R_{M})$, and denote it by $\mathrm{Mod}(R_{M})$, viewed as an $\infty$-category (or more concretely, a dg-category).
In the case of $\mathbb{G}\neq\mathbb{R}$, to compare with sheaf theory, we introduce a noncommutative algebra as follows. We first introduce the following quiver: The set of vertices is $\mathbb{R}/\mathbb{G}$. For $[c],[c^{\prime}]\in\mathbb{R}/\mathbb{G}$, the set of arrows from $[c]$ to $[c^{\prime}]$ is identified with the set
$\left\{d\in\mathbb{R}_{\geq 0}\mathrel{}\middle|\mathrel{}[c+d]=[c^{\prime}]\right\}.$ (2.3)
We denote the morphism corresponding to $d$ by $e_{[c],d}$. We denote the associated quiver algebra by $\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]$. An element of this algebra is of the form
$\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}$ (2.4)
where $a_{[c],d}\in\mathbb{K}$ are zero except for finitely many $d$ for each $[c]$. For two elements, $\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d},\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}e_{[c],d}$, we define the product by
$\left(\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}e_{[c],d}\right)\cdot\left(\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}\right)=\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d^{\prime\prime}\in\mathbb{R}_{\geq 0}}\left(\sum_{d+d^{\prime}=d^{\prime\prime}}b_{[c+d],d^{\prime}}a_{[c],d}\right)e_{[c],d^{\prime\prime}}$ (2.5)
@@
For each $c$, the identity morphism is denoted by $e_{c}\in\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]$. For $\ell>0$, we denote the ideal generated by the arrows represented by the positive numbers greater than $\ell$ by $\mathfrak{m}(\ell)$. We take the completion of the algebra with $\mathfrak{m}(\ell)$-adic topology and denote it by $L_{0}^{\mathbb{G}}$. More explicitly, an element of $L_{0}^{\mathbb{G}}$ is of the form
$\prod_{[c]\in\mathbb{R}/\mathbb{G}}\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}e_{[c],d}$ (2.6)
@@
Let $M$ be a finite-dimensional persistence module, namely, a functor from the poset category $(\mathbb{R},\geq)$ to the category of finite-dimensional vector spaces. For each $c\in\mathbb{R}$, we denote the image under the functor by $M_{c}$. We suppose that there exists $L\in\mathbb{R}$ such that $M_{c}=0$ for any $c<L$. In the following, we will see that $\prod_{c\in\mathbb{R}}M_{c}$ carries a $L_{0}^{\mathbb{O}}$-module structure.
For $c\leq c^{\prime}$, we denote the structure morphism by $t_{c,c^{\prime}}\colon M_{c}\rightarrow M_{c^{\prime}}$. We set $N_{-c}:=M_{c}$. Take an element $\prod_{c\in\mathbb{R}}n_{c}=\prod_{c>-L}n_{c}\in\prod_{c\in\mathbb{R}}N_{c}$. We also set $e_{[c],d}=:f_{c,c+d}$. The action is defined by
@@
On $e_{[c],d}$, $T^{d^{\prime}}\in\Lambda_{0}$ acts on it by $e_{[c],d}\mapsto e_{[c],d+d^{\prime}}$. This defines the desired action. ∎
@@ 2.3 Real valuation
Let $A$ be an integral domain with a map $v\colon A\rightarrow\mathbb{R}_{\geq 0}$. We say $(A,\mathfrak{v})$ is a real valuation ring if
@@
$v(x):=\min\left\{c\in\mathbb{R}_{\geq 0}\mathrel{}\middle|\mathrel{}a\in T^{c}\mathfrak{m}\right\}.$ (2.10)
This obviously gives a real valuation of $\Lambda_{0}$.

### v3

@@ 1 Introduction
1.1 Main result
In symplectic geometry, there exists a series of extra variables, which are (expected to be) just various aspects of the same variable.
@@ 1.1 Main result
    In microlocal sheaf theory, it is called $t$ and introduced by Tamarkin [Tam18] to study sheaf quantizations, which was also envisioned by Sato in relation with WKB analysis [SK14].
@@
For example, in [Kuw24], we discussed the relationship between (1) and (2); By using exact WKB analysis for $\hbar$-differential operators, we construct sheaf quantization. In the process, the value in (2) is identified with (1). Also, some hints of the relationship between (1) and (3) can be seen from the work of Tamarkin and [AI20, IK]; In [AI20], the relation between symplectic energy and $t$ is explained, which is the same role as the Novikov variable. In [IK], this point of view is further expanded to develop a sheaf-theoretic construction parallel to Lagrangian Floer theory.
@@
We would like to formulate our main theorem now. Let $\mathbb{K}$ be a field. Let $X$ be a manifold and $\mathbb{R}_{t}$ be the real line. We denote the discrete additive group of the real numbers by $\mathbb{R}_{d}$. We then have the equivariant derived category $\mathrm{Sh}^{\mathbb{R}_{d}}(X\times\mathbb{R}_{t},\mathbb{K})$ of $\mathbb{K}$-modules sheaves. We quotient this category by the sheaves with non-positive microsupport and denote it by $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$. The non-equivariant version of this category was originally introduced by Tamarkin [Tam18], and the equivariant version was later introduced by the present author [Kuw24]. On the other hand, we have the universal Novikov ring $\Lambda_{0}$ over $\mathbb{K}$. We denote the derived category of $\Lambda_{0}$-module sheaves by $\mathrm{Sh}(X,\Lambda_{0})$.
@@
$\mathfrak{A}\colon\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(X,\Lambda_{0}).$ (1.1)
The almost images are sheaves valued in derived complete modules over $\Lambda_{0}$.
Here the term “almost” comes from the almost mathematics [GR03]. Roughly speaking, the above theorem without “almost” holds after negating almost zero modules. The precise meaning will be explained in the body of the paper. The use of the almost mathematics in symplectic geometry also happened in [Fuk].
@@
Remark 1.2.
Here let us introduce some previous works.
  1. 1.
    In [Kuw24], we found the Novikov ring a posteriori after the construction of the equivariant version of Tamarkin’s category. Namely, we showed that the said category is linear over the Novikov ring. In this paper,we show that Novikov sheaves and the equivariant Tamarkin category is actually (almost) equivalent, which is stronger that the former one. In particular, this equivalence allows us to have an alternative model for the category of sheaf quantizations.
  2. 2.
    The holonomic Riemann–Hilbert correspondence by D’Agnolo–Kashiwara [DK16] is formulated and proved by using a variant of Tamarkin category, called enhanced ind-sheaves. In [Kuw21], we use the sheaves valued in modules over the finite Novikov ring $\mathbb{K}[\mathbb{R}_{\geq 0}]$ to reformulate the holonomic Riemann–Hilbert correspondence. One can reprove this result by using this paper with some understanding on the relationship between enhanced ind-sheaves and Tamarkin category. This will be further explained in [KK26].
  3. 3.
    After completing this manuscript, Bingyu Zhang informed the author that Vaintrob sketched a similar result for non-negatively microsupported sheaves in his unpublished manuscript [Vai17].
1.2 Why is the theorem important?
In this section, we explain the reason why the stated main theorem seems to be important.
1.2.1 Symplectic geometry
The original non-equivariant version of Tamarkin’s category [Tam18] is known to capture many symplectic geometric information of $T^{*}X$ and its Lagrangian submanifolds. Viterbo [Vit] envisioned that the Tamarkin category contains filtered Fukaya category of exact Lagrangians of $T^{*}X$, and constructed a baby version of the functor (a sheaf-theoretic construction of the expected images were constructed by Guillermou [Gui16]). The idea is very canonical: Consider the family Floer functor with respect to the cotangent fibers, then one gets a sheaf on $X$. Since the Floer cohomology is filtered in the exact case, one can translate it to the additional $\mathbb{R}$-direction of the Tamarkin category.
If one expects a generalization of the family Floer functor to the non-exact case, the functor should take its value in the category of the sheaves over the Novikov ring, since the Floer theory is defined over the Novikov ring [FOOO09], due to the infiniteness of the number of holomorphic disks.
On the other hand, in the non-exact case, the works of [Kuw24, IK] show that the equivariant Tamarkin category works well to treat non-exact Lagrangians. So, in particular, [IK] conjectures that a version of non-exact Fukaya category of $T^{*}X$ is embedded to $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$. An integral version of the conjecture is proved in  [KPS] without using family Floer theory, but using [GPS24].
To prove this conjecture by using the family Floer theory, one has to relate $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$ and the category of sheaves over the Novikov ring, which is the natural target of the family Floer functor. The main theorem explains the relationship precisely.
Aside from this application, we will explain two applications of (the philosophy) of the theorem.
The first one is a first step toward nonconic microlocal sheaf theory of sheaves valued in modules over real valuation rings. Namely, for a manifold $X$, a real valuation ring $R$, and a sheaf $\mathcal{E}$ of modules over $R$, we can define a subset $\mu\mathrm{supp}(\mathcal{E})\subset T^{*}M$ by interpreting Tamarkin’s non-conic microsupport. This refines Kashiwara–Schapira’s microsupport [KS94].
@@ 1.2.1 Symplectic geometry
1.2.2 Algebraic asymptotic analysis
We also discuss an application to asymptotic analysis. Let $f(\hbar)$ be a function which is asymptotically expandable toward $\hbar\to 0$ (say, from the real positive direction), namely it has an associated power series expansion. In this expansion, functions like $e^{-c/\hbar}$ ($c\in\mathbb{R}_{>0}$) are ignored, since they have zero asymptotic expansion. However, sometimes, one can expand the function further into a formal power series of $e^{-c/\hbar}$, by seeing more precise asymptotics. The resulting expansion is called transseries, and by interpreting $e^{-c/\hbar}$ as $T^{c}$, the ring of transseries is a Novikov ring.
On the other hand, in [Kuw24], the author constructed an object of $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$ from a second-order $\hbar$-differential equation over a Riemann surface, by using exact WKB analysis: From exact WKB analysis, one obtains a global version of Stokes matrices (called Voros matrices). Gluing by Voros matrices give an object of $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$.
The higher-order version of exact WKB analysis is not well-established yet, it has been expected that the resulting solutions and Stokes matrices are expressed by transseries. So, the natural place where we can find solutions would be sheaves over the Novikov ring. Then, to generalize the picture from [Kuw24] to the higher-order cases, we again need to relate $\mathrm{Sh}^{\mathbb{R}_{d}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K})$ and the category of sheaves over the Novikov ring, which is again the job of our main theorem in this paper.
@@ 1.2.2 Algebraic asymptotic analysis
Let $\mathbb{K}$ be a field. All the operations in this paper should be understood in the derived sense unless otherwise stated. For a $\mathbb{K}$-algebra $A$, $\mathrm{Mod}(A)$ (resp. $\mathrm{Mod}^{\heartsuit}(A)$) is the derived $\infty$-category of right $A$-modules (resp. the abelian category of right $A$-modules).
@@ Acknowledgment
This work is supported by JSPS KAKENHI Grant Numbers 22K13912, 23H01068, 20H01794, and 25K21660. I’d like to thank Yuichi Ike for related discussions. I also thank Tomohiro Asano, Takumi Arai, Bingyu Zhang, and anonymous referees for sevreal helpful comments on earlier drafts.
@@ 2.1 Definition
Let $\mathbb{R}$ be the 1-dimensional Euclidean vector space. Let $\mathbb{G}$ be a subgroup of $\mathbb{R}$. Then $\mathbb{R}_{\geq 0}\cap\mathbb{G}$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$. We denote the indeterminate corresponding to $a\in\mathbb{R}_{\geq 0}\cap\mathbb{G}$ by $T^{a}$. Let $|\cdot|$ be the Euclidean norm of $\mathbb{R}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]$ generated by $T^{a}$’s with $a>r$ by $\mathfrak{m}(r)$. Obviously, $\mathfrak{m}(r^{\prime})\supset\mathfrak{m}(r)$ if $r>r^{\prime}$. Hence $\mathbb{K}[\mathbb{R}_{\geq 0}\cap\mathbb{G}]/\mathfrak{m}(r)$ forms a projective system.
@@
$\Lambda_{0}:=\Lambda_{0}^{\mathbb{R}}\cong\lim_{\begin{subarray}{c}\longleftarrow\\ a\rightarrow+\infty\end{subarray}}\mathbb{K}[\mathbb{R}_{\geq 0}]/T^{a}\mathbb{K}[\mathbb{R}_{\geq 0}].$ (2.2)
@@ 2.2 Modules over the Novikov ring
Let $X$ be a topological space. Let $R_{X}$ be the constant sheaf valued in $R$. We denote the abelian category $R_{X}$-modules by $\mathrm{Sh}^{\heartsuit}(X,R)$. We similarly consider the derived category of $\mathrm{Sh}^{\heartsuit}(X,R)$, and denote it by $\mathrm{Sh}(X,R)$, viewed as an $\infty$-category (or more concretely, a dg-category). Note that $\mathrm{Sh}(X,R)$ is equivalent to the $\infty$-category of sheaves valued in $\mathrm{Mod}(R)$ if $X$ is hypercomplete [Sch].
In the case of $\mathbb{G}\neq\mathbb{R}$, to compare with sheaf theory, we introduce a noncommutative algebra as follows. We first introduce the following quiver: The set of vertices is $\mathbb{R}/\mathbb{G}$. For $c],[c^{\prime}]\in\mathbb{R}/\mathbb{G}$, the set of arrows from $[c]$ to $c^{\prime}]$ is defined as
$\left\{T_{[c]}^{d}\mathrel{}\middle|\mathrel{}d\in\mathbb{G}\cap\mathbb{R}_{\geq 0}\right\}$ (2.3)
We denote the associated quiver algebra by $\mathbb{K}[Q(\mathbb{R}/\mathbb{G})]$. An element of this algebra is of the form
$\left(\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}}$ (2.4)
where $a_{[c],d}\in\mathbb{K}$ are zero except for finitely many $d$ for each $[c]$. For two elements, $\left(\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}},\left(\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}}$, we define the product by
$\left(\sum_{d\in\mathbb{R}_{\geq 0}}b_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}}\cdot\left(\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}}=\left(\sum_{d^{\prime\prime}\in\mathbb{R}_{\geq 0}}\left(\sum_{d+d^{\prime}=d^{\prime\prime}}b_{[c+d],d^{\prime}}a_{[c],d}\right)T_{[c]}^{d^{\prime\prime}}\right)_{[c]\in\mathbb{R}/\mathbb{G}}$ (2.5)
@@
For $\ell>0$, we denote the ideal generated by the arrows represented by the positive numbers greater than $\ell$ by $\mathfrak{m}(\ell)$. Explicitly,
$\mathfrak{m}(\ell)=\left\{\left(\sum_{d\in\mathbb{G}\cap\mathbb{R}_{\geq 0}}a_{[c],d}T^{d}_{[c]}\right)_{[c]\in\mathbb{R}/\mathbb{G}}\mathrel{}\middle|\mathrel{}a_{[c],d}=0\text{ for }d\leq\ell\right\}$ (2.6)
We take the completion of the algebra with $\mathfrak{m}(\ell)$-adic topology and denote it by $L_{0}^{\mathbb{G}}$. More explicitly, an element of $L_{0}^{\mathbb{G}}$ is of the form
$\left(\sum_{d\in\mathbb{R}_{\geq 0}}a_{[c],d}T_{[c]}^{d}\right)_{[c]\in\mathbb{R}/\mathbb{G}}$ (2.7)
@@
Let $M$ be a persistence module, namely, a functor from the poset category $(\mathbb{R},\geq)$ to the category of vector spaces. For each $c\in\mathbb{R}$, we denote the image under the functor by $M_{c}$. We suppose that there exists $L\in\mathbb{R}$ such that $M_{c}=0$ for any $c<L$. In the following, we will see that $\prod_{c\in\mathbb{R}}M_{c}$ carries a $L_{0}^{\mathbb{O}}$-module structure.
For $c\leq c^{\prime}$, we denote the structure morphism by $t_{c,c^{\prime}}\colon M_{c}\rightarrow M_{c^{\prime}}$. We set $N_{-c}:=M_{c}$. Take an element $\prod_{c\in\mathbb{R}}n_{c}=\prod_{c\leq-L}n_{c}\in\prod_{c\in\mathbb{R}}N_{c}$. We also set $T_{[c]}^{d}=:f_{c,c+d}$. The action is defined by
@@
On $T_{[c]}^{d}$, the element $T^{d^{\prime}}\in\Lambda_{0}$ acts on it by $T_{[c]}^{d}\mapsto T_{[c]}^{d+d^{\prime}}$. This defines the desired action. ∎
@@ 2.3 Real valuation
Let $A$ be an integral domain with a map $v\colon A\rightarrow\mathbb{R}_{\geq 0}\cup\left\{+\infty\right\}$. We say $(A,v)$ is a real valuation ring if
@@
$v(a):=\sup\left\{c\in\mathbb{R}_{\geq 0}\mathrel{}\middle|\mathrel{}a\in T^{c}\mathfrak{m}\right\}.$ (2.11)
where $\mathfrak{m}$ is the unique maximal ideal of $\Lambda_{0}$. This obviously gives a real valuation of $\Lambda_{0}$.